OpenStudy (anonymous):

please help this one is really hard for me Simplify f(x)=2x^2+3x+1/2x+1 Using complete sentences, explain why f(1) =2, f(0)=1 and f(-1)=0, yet f(-1/2) is undefined. Be sure to show your work

geerky42 (geerky42):

Because f(-1/2) has zero in denominator.

OpenStudy (anonymous):

Factor the top part of the function and simplify

geerky42 (geerky42):

In denominator, 2x + 1 = 0 2x = -1 x = -1/2 f(x) is undefined when x = -1/2

OpenStudy (tkhunny):

It is likely to be more clear if you write it correctly. f(x)=(2x^2+3x+1)/(2x+1) The parentheses are NOT optional.